Let n be the unknown number of customers that visit a store on the day of a sale. The number of customers that make a purchase is Y | n ~ Binomial(n, theta) where theta is the known probability of making a purchase given the customer visited the store. The prior is n ~ Poisson(5). Assuming theta is known and n is the known parameter, plot the posterior distribution of n for all combinations of Y=0,5,10 and theta=0.2,0.5 and comment on the effect of Y and theta on the posterior.
Answer #1

hl 2 ) 2 od (5つ) n-2 セフ): beremes sum b exponent 4-1 hnt NIY ~ Poisson (5-50丿 Poisson (4 ) for Y:0, wen 冫0.2

above is the required theory for this problem.

following is the required code in R to plot the posterior densities with varying theta and y:

for(theta in c(0.2,0.5)) 14 15 16 17 18 19 20 21 for(k in c(0,5,10)) x<-k:50 y<-exp(- (5-5*theta)) C(5-5*theta)A(x-k))/factorial(x-k) plot (x , y , xl í m=c(0,50).ylim=c(0,1),x1ab-k , main-theta,ylab=posterior density)following are the plots when

theta=0.2(x labels are the y values, header in the theta value)

0.2 0 0 10 20 30 40 50 0

0.2 0 10 20 30 40 50 5

0.2 0 10 20 30 40 50 10

from these above plots(when theta=0.2) with increasing y values the posterior distribution shifts to the right.

when theta=0.5(x labels are the y values, header in the theta value)

posterior density 0.0 04 08 8 8posterior density 0.0 04 08

0.5 0 10 20 30 40 50 10here also with increasing y the posterior distribution shifts to the right.

it is to be noted that as theta increases mode also increases for the posterior density.